### v3

# TropNNC: Structured Neural Network Compression Using Tropical Geometry

Abstract
We present TropNNC, a framework for compressing neural networks with linear and convolutional layers and ReLU activations using tropical geometry. By representing a network’s output as a tropical rational function, TropNNC enables structured compression via reduction of the corresponding tropical polynomials. Our method refines the geometric approximation of previous work by adaptively selecting the weights of retained neurons. Key contributions include the first application of tropical geometry to convolutional layers and the tightest known theoretical compression bound. TropNNC requires only access to network weights – no training data – and achieves competitive performance on MNIST, CIFAR, and ImageNet, matching strong baselines such as ThiNet and CUP.
11footnotetext: Corresponding author.
1 Introduction
Deploying deep neural networks on resource-constrained devices remains challenging due to their substantial computational and storage demands. Initial attempts to reduce network complexity include unstructured pruning. which alters the network’s structure by eliminating individual weights. While effective in reducing network size, unstructured pruning presents challenges in practical applications (Wen et al., 2016). To overcome its limitations, structured pruning methods, such as channel-level pruning, have been proposed (He and Xiao, 2024). Notably, the ThiNet framework by Luo et al. (2017) prunes entire filters or channels, maintaining the network’s original structure. This structured approach ensures compatibility with existing deep learning libraries and offers several advantages: it significantly reduces memory footprint, and facilitates further compression and acceleration through methods like parameter quantization.
Parallel to these advancements, tropical geometry (Maclagan and Sturmfels, 2021) has emerged as a promising mathematical framework with applications in machine learning (Maragos et al., 2021; Gärtner and Jaggi, 2008) and beyond. Recently, it has been applied to the theoretical study of neural networks. For example, Zhang et al. (2018) demonstrated the equivalence of ReLU-activated neural networks with tropical rational mappings. Together with other works, like those of Charisopoulos and Maragos (2018); Alfarra et al. (2023), they used tropical geometry to compute bounds on the number of linear regions of neural networks equal to the one in (Montufar et al., 2014). Smyrnis and Maragos (2020) use tropical geometry to develop pruning methods.
Contributions.
In this paper, we explore the application of tropical geometry in the compression of ReLU neural networks. Our contributions include:
  • Proposing TropNNC, an algorithm that leverages tropical geometry and the Hausdorff distance for the structured compression of neural networks. TropNNC compresses convolutional networks layer-wise by approximating the zonotopes corresponding to each layer using an iterative process. Our algorithm is data agnostic: it does not require a training dataset.
  • Refining the bound for the functional approximation of tropical polynomials presented in (Misiakos et al., 2022), and using it to provide stronger theoretical compression guarantees.
  • Evaluating our algorithm empirically on MNIST, CIFAR and ImageNet datasets. Our method outperforms prior tropical geometrical pruning methods, and achieves competitive performance compared to the data-driven ThiNet, and superior performance compared to CUP (Duggal et al., 2021), particularly in the VGG architecture.
This work demonstrates the potential of tropical geometry in enhancing neural network compression techniques. To the best of our knowledge, it is the first tropical geometric pruning method that compresses convolutional layers and is competitive with strong baselines. Even if currently not state-of-the-art, we believe that our contribution is a necessary step towards this direction. Related work is provided after the definition of our algorithm to enable detailed comparisons. Proofs of theoretical results are provided in the appendix.
2 Tropical Geometry of Neural Networks
Tropical algebra studies matrix-vector operations based on the arithmetic of the tropical semiring Cuninghame-Green (1979); Butkovič (2010). Tropical geometry is the counterpart of algebraic geometry in the tropical setting. The tropical semiring can refer to either the min-plus semiring or the max-plus semiring. In this work, we adhere to the convention of using the max-plus semiring $(\mathbb{R}_{\mathrm{max}},\vee,+)$, defined as the set $\mathbb{R}_{\mathrm{max}}=\mathbb{R}\cup\{-\infty\}$ equipped with two binary operations: $\vee$ (ordinary max) and ++ (ordinary sum).
Within the max-plus semiring, we can define tropical polynomials, which correspond to convex piecewise linear functions, and tropical rational functions, which correspond to general continuous piecewise linear functions. We can also define Newton polytopes of tropical polynomials, which connect tropical algebra with polytope theory. For a detailed introduction to these concepts refer to the appendix.
2.1 Neural Networks with Piecewise Linear Activations
Tropical geometry provides a mathematical framework for analyzing neural networks with piecewise linear activation functions. In this work, we focus on ReLU-activated networks.
ReLU Activations.
Consider a network which consists of an input layer $\mathbf{x}=(x_{1},\ldots,x_{d})$, a hidden layer $\mathbf{f}=(f_{1},\ldots,f_{n})$ of ReLU units, and an output layer $\mathbf{v}=(v_{1},\ldots,v_{m})$. The input, hidden, and output layers are connected through linear transformations represented by matrices $\mathbf{A}$ and $\mathbf{C}$. Each neuron $i$ has input weights and bias given by $\mathbf{A}_{i,:}=(\mathbf{a}_{i}^{T},b_{i})$ and output weights $\mathbf{C}_{:,i}^{T}=(c_{1i},\ldots,c_{mi})$. We assume the output layer has no bias. Such a network is depicted in Figure 1.
Figure 1: Neural network with one hidden ReLU layer. The first linear layer has weights $\{\mathbf{a}_{i}^{T}\}$ with bias $\{b_{i}\}$ corresponding to node $i\in[n]$ and the second has weights $\{c_{ji}\}$ between nodes $j\in[m],i\in[n]$
The output of the ReLU unit $i$ is given by:
$f_{i}(\mathbf{x})=\mathrm{ReLU}(\mathbf{a}_{i}^{T}\mathbf{x}+b_{i})=\max\{\mathbf{a}_{i}^{T}\mathbf{x}+b_{i},0\}.$
This expression represents a tropical polynomial of rank 2, with one term being the constant 0. The extended Newton polytope $\mathrm{ENewt}(f_{i})$ of $f_{i}$ is an edge with one endpoint at the origin $\mathbf{0}$ and the other endpoint at $(\mathbf{a}_{i}^{T},b_{i})$. The $j$-th component of the output layer $v_{j}$ can be computed as follows:
$v_{j}=\sum_{i\in[n]}c_{ji}f_{i}=\sum_{i:c_{ji}>0}|c_{ji}|f_{i}-\sum_{i:c_{ji}<0}|c_{ji}|f_{i}=p_{j}-q_{j}.$
In the above expression, $|c_{ji}|f_{i}$ are tropical polynomials. Thus, $p_{j}$ and $q_{j}$ are tropical polynomials formed by the addition of tropical polynomials. Consequently, $v_{j}$ is a tropical rational function. We call $p_{j}$ the positive polynomial and $q_{j}$ the negative polynomial of $v_{j}$.
Zonotopes.
The extended Newton polytope of $|c_{ji}|f_{i}$ is an edge with one endpoint at the origin $\mathbf{0}$ and the other at $|c_{ji}|(\mathbf{a}_{i}^{T},b_{i})$. The extended Newton polytope $P_{j}$ of $p_{j}$ is the Minkowski sum of the positive generators $\{|c_{ji}|(\mathbf{a}_{i}^{T},b_{i}):c_{ji}>0\}$, and the polytope $Q_{j}$ of $q_{j}$ is the Minkowski sum of the negative generators $\{|c_{ji}|(\mathbf{a}_{i}^{T},b_{i}):c_{ji}<0\}$. Thus, $P_{j},Q_{j}$ are zonotopes. We refer to $P_{j}$ as the positive zonotope and $Q_{j}$ as the negative zonotope of $v_{j}$.
3 Approximation based on Hausdorff distance
In this section, we present our refined theorem, which uses the Hausdorff distance in its standard continuous form to bound the error between two tropical polynomials. In the following, the distance between two points $\mathbf{u}$ and $\mathbf{v}$ is denoted as $\text{dist}(\mathbf{u},\mathbf{v}):=\|\mathbf{u}-\mathbf{v}\|$, where $\|\cdot\|$ denotes the standard $L^{2}$ Euclidean norm. The distance between a point $\mathbf{u}$ and a set $V$ is defined as $\operatorname{dist}(\mathbf{u},V)=\operatorname{dist}(V,\mathbf{u}):=\inf_{\mathbf{v}\in V}\|\mathbf{u}-\mathbf{v}\|$.
Definition 1 (Hausdorff distance).
Let $S,\tilde{S}$ be two subsets of $\mathbb{R}^{d}$. The Hausdorff distance H(S,S~)H($\tilde{S}$) of the two sets is defined as
$H(S,\tilde{S}):=\max\left\{\sup_{\mathbf{u}\in S}\mathrm{dist}(\mathbf{u},\tilde{S}),\sup_{\mathbf{v}\in\tilde{S}}\mathrm{dist}(S,\mathbf{v})\right\}$
In the case of polytopes $P,\tilde{P}$, due to their convexity and compactness, the suprema in the above expression are attained, and in fact by points in the vertex sets $V_{P},V_{\tilde{P}}$ of the polytopes.
[…]
Theorem 2.
[…]
where $B=\{\mathbf{x}\in\mathbb{R}^{d}:\|\mathbf{x}\|\leq r\}$ and $\rho=\sqrt{r^{2}+1}$.
Since $V_{P}\subseteq P,V_{\tilde{P}}\subseteq\tilde{P}$, we have that $\mathrm{dist}(\mathbf{u},\tilde{P})\leq\mathrm{dist}(\mathbf{u},V_{\tilde{P}}),\mathrm{dist}(P,\mathbf{v})\leq\mathrm{dist}(V_{P},\mathbf{v})$. The equality in the aforementioned inequalities is achieved only in special cases and thus the bound H(P,P~)H($\tilde{P}$) we provide is in general tighter than DH(P,P~)DH($\tilde{P}$) from Misiakos et al. (2022).
By applying the triangle inequality and Theorem 2 we obtain the following result for two neural networks of compatible input and output dimension.
Corollary 3.
Let $\mathbf{v}$ and $\tilde{\mathbf{v}}$ be the outputs of two neural networks as in Figure 1. Then, the following inequality holds:
$\frac{1}{\rho}\max_{\mathbf{x}\in B}\|\mathbf{v}(\mathbf{x})-\tilde{\mathbf{v}}(\mathbf{x})\|_{1}\leq\sum_{j=1}^{m}\left(H(P_{j},\tilde{P}_{j})+H(Q_{j},\tilde{Q}_{j})\right),$
where $\|\cdot\|_{1}$ denotes the $L^{1}$ norm.
4 Compression Algorithm
From Cor. 3 it is evident that to compress a ReLU network consisting of a pair of consecutive linear layers, one has to choose compressed weight matrices $\widetilde{\mathbf{A}}\in\mathbb{R}^{K\times(d+1)}$ and $\widetilde{\mathbf{C}}\in\mathbb{R}^{m\times K}$ with $K<n$ such that $\widetilde{P}_{j}\approx P_{j}$ and $\widetilde{Q}_{j}\approx Q_{j}$ for all $j\in[m]$, where the approximation relation is in terms of the Hausdorff distance between the zonotopes.
4.1 Single output
If we have $m=1$ then the problem can readily be translated into a zonotope approximation problem—the task of approximating a zonotope with another zonotope that has fewer generators. This problem is known as zonotope order reduction (Yang and Scott, 2018). In our case, the approximation must happen in terms of the Hausdorff distance.
Algorithm for single output network.
For single-output networks, our approach uses K-Means to cluster positive and negative generators, and replaces each cluster with a single representative. Unlike Misiakos et al. (2022), who define the representative as the center/mean of the cluster, we take the representative to be the sum of the cluster.
Our Algorithm 1 is depicted below. Figure 2 illustrates an example execution of TropNNC compared to the method of Misiakos et al. (2022). For our algorithm, the following bound and its corollary hold.
Algorithm 1 TropNNC, Single output
1: Split the generators $|c_{i}|(\mathbf{a}_{i}^{T},b_{i})$ into positive and negative generators:
$\{$ $|c_{i}|(\mathbf{a}_{i}^{T},b_{i}):c_{i}>0\},$
$\{$ $|c_{i}|(\mathbf{a}_{i}^{T},b_{i}):c_{i}<0\}.$
2: Execute K-means with $K/2$ centers on the positive generators $\{|c_{i}|(\mathbf{a}_{i}^{T},b_{i}):c_{i}>0\}$, and $K/2$ centers on the negative generators $\{|c_{i}|(\mathbf{a}_{i}^{T},b_{i}):c_{i}<0\}$.
3: Obtain positive and negative cluster representatives:
$\{$ $|\tilde{c}_{i}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i}):i\in C_{+}\},$
$\{$ $|\tilde{c}_{i}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i}):i\in C_{-}\},$
where $C_{+}\sqcup C_{-}=[K]$ and $|\tilde{c}_{i}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i})$ is the sum of the generators of cluster $i$.
4: For each $i\in[K]$ construct a (hidden layer) neuron with input weights and bias $|\tilde{c}_{i}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i})$.
5: For each constructed neuron $i$ set the output weight to 11 if the neuron corresponds to a representative of a positive cluster ($i\in C_{+}$), otherwise set it to -1.
(a) Original
(b) Misiakos et al.
(c) TropNNC
Figure 2: Example execution of TropNNC compared to the method of Misiakos et al. (2022).
Proposition 4.
Suppose the clusters K are enough so that for every cluster, no 2 generators of the cluster form an obtuse angle. Then, single-output TropNNC produces a neural network with output $\tilde{v}$ satisfying:
[…]
where $\delta_{max}$ is the largest distance from a point to its corresponding cluster center.
Corollary 5.
The above bound is tighter than the bound of Zonotope K-means of Misiakos et al. (2022).
4.2 Multi-output
We now consider the multi-output case with $m\in\mathbb{N}$. Notice an interesting property of the zonotopes $P_{j},Q_{j}$: they share the directions $(\mathbf{a}_{i}^{T},b_{i})$ of their generators. For instance, output $v_{1}$ might have a positive generator $|c_{1i}|(\mathbf{a}_{i}^{T},b_{i})$ of zonotope $P_{1}$, while output $v_{2}$ might have a negative generator $|c_{2i}|(\mathbf{a}_{i}^{T},b_{i})$ of zonotope $Q_{2}$. These generators are parallel to each other, with common direction $(\mathbf{a}_{i}^{T},b_{i})$. Hence, our original positive and negative zonotopes have parallel generators, and we aim to approximate them with new positive and negative zonotopes, also with parallel generators. We refer to this complex approximation problem as simultaneous zonotope approximation.
Example 1.
[…]
(a) Zonotope of $1^{st}$ output
(b) Zonotope of $2^{nd}$ output
(c) Sim. zonotope approximation
Figure 3: Example of simultaneous zonotope approximation for a network with 2 outputs and 2 hidden neurons
4.2.1 Non-iterative Algorithm for multi-output network.
For the multi-output case, we perform clustering of similar neurons based on the set of clustering vectors $\left\{(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T}),i\in[n]\right\}$, and replace each cluster with a single representative. Unlike Misiakos et al. (2022), who take the cluster center/mean as the representative, we form the representative as follows: For every cluster $k\in[K]$ with clustered neuron indexes $I_{k}$ and vectors $\left\{(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T}),i\in I_{k}\right\}$ take $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ to be the mean of $\left\{(\mathbf{a}_{i}^{T},b_{i}),i\in I_{k}\right\}$, take $\widetilde{\mathbf{C}}_{:,k}^{T}$ to be the sum of $\left\{\mathbf{C}_{:,i}^{T},i\in I_{k}\right\}$, and form the representative of the cluster $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k},\widetilde{\mathbf{C}}_{:,k}^{T})$. The complete procedure is shown in Algorithm 2.
Algorithm 2 (Non-iterative) TropNNC for Multi-output networks
1: Execute clustering on the clustering vectors $(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T})$ for $i\in[n]$, forming $K$ clusters $\{(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T})\mid i\in I_{k}\}$ for $k\in[K]$.
2:For each $k\in[K]$, form the cluster representative $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k},\widetilde{\mathbf{C}}_{:,k}^{T})$ as follows:
  • (i)
    Compute $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ as the mean of the input weights and biases of the vectors in the cluster, i.e., the mean of the set $\{(\mathbf{a}_{i}^{T},b_{i})\mid i\in I_{k}\}$.
  • (ii)
    Compute $\widetilde{\mathbf{C}}_{:,k}^{T}$ as the sum of the output weights of the vectors in the cluster, i.e., the sum of the set $\{\mathbf{C}_{:,i}^{T}\mid i\in I_{k}\}$.
3: Construct the new hidden layer:
  • (i)
    For the input weights, set the $k$-th row of the weight-bias matrix to $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$.
  • (ii)
    For the output weights, set the $k$-th column to $\widetilde{\mathbf{C}}_{:,k}$.
4.2.2 Iterative Algorithm for multi-output network.
To improve the approximation of Algorithm 2, we formulate an optimization problem that takes the output of Algorithm 2 and with an iterative process produces weights that achieve a better simultaneous zonotope approximation.
Motivated by Algorithm 1, assuming the number of null neurons are few (see appendix), we want in terms of every output $j$ the cluster representative $\widetilde{C}_{j,k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ to be a close approximation to the cluster sum $\sum_{i\in I_{k}}C_{j,i}(\mathbf{a}_{i}^{T},b_{i})$. Thus, for every cluster $k$, we have unknowns $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k},\widetilde{\mathbf{C}}_{:,k}^{T})=(\tilde{a}_{k1},\ldots,\tilde{a}_{kd},\tilde{b}_{k},\widetilde{C}_{1k},\ldots)$, and we wish to find a solution which minimizes the following criterion:
$\sum_{j=1}^{m}\left\|\tilde{C}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})-\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\right\|^{2},$ (1)
where $m$ is the number of outputs, and $I_{k}$ is the set of neurons of cluster $k$.
The above optimization problem can be solved by means of iterative alternating minimization. Specifically:
1. Fixing the input weights $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ and minimizing with respect to $\widetilde{\mathbf{C}}_{:,k}$, the terms of the sum are independent and thus can be minimized separately. For each term of the sum, the optimal error occurs if we project the sum $\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})$ onto $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$. We have:
$\widetilde{C}_{jk}=\frac{\left\langle\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i}),(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\rangle}{\left\|(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|^{2}}$ (2)
2. Fixing the output weights $\widetilde{\mathbf{C}}_{:,k}$, and minimizing with respect to the input weights $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$, we take the derivative of criterion (1) with respect to $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ and set it to zero. We have:
$2\sum_{j=1}^{m}\left[\widetilde{C}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})-\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\right]\widetilde{C}_{jk}=0\Leftrightarrow$
$\Leftrightarrow(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\sum_{j=1}^{m}\widetilde{C}_{jk}^{2}=\sum_{j=1}^{m}\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\widetilde{C}_{jk}\Leftrightarrow$
$(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})=\frac{\sum_{j=1}^{m}\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\widetilde{C}_{jk}}{\sum_{j=1}^{m}\widetilde{C}_{jk}^{2}}$ (3)
We can initialize the iteration with the representative obtained from Algorithm 2. The resulting procedure is detailed in Algorithm 3. We note that, in practice, the number of null generators is not negligible, and thus criterion (1) also constitutes a heuristic method. For this reason, the number of iterations should not be excessive. We provide Proposition 6 for the approximation error of our algorithm.
Algorithm 3 Iterative TropNNC for Multi-output networks
1: Form the initial cluster representatives $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k},\widetilde{\mathbf{C}}_{:,k}^{T})$ by executing the non-iterative algorithm.
2:for $\text{iter}=1\text{ to }num\_iter$ do
3:  for $k\in[K]$ do
4:   for $j\in[m]$ do
5:    $\widetilde{C}_{jk}\leftarrow\frac{\langle\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i}),(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\rangle}{\|(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\|^{2}}$
6:   end for
7:   $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\leftarrow\frac{\sum_{j=1}^{m}\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\widetilde{C}_{jk}}{\sum_{j=1}^{m}\widetilde{C}_{jk}^{2}}$
8:  end for
9:end for
10: Construct the new linear layer as in the non-iterative algorithm.
Proposition 6.
Suppose the clusters K are enough so that for every cluster, no two $(\mathbf{a}_{i}^{T},b_{i}),(\mathbf{a}_{i^{\prime}}^{T},b_{i^{\prime}})$ of the cluster form an obtuse angle. Then, a variant of Iterative TropNNC produces a neural network with output $\tilde{v}$ satisfying:
$\frac{1}{\rho}\max_{\mathbf{x}\in B}\|\mathbf{v}(\mathbf{x})-\tilde{\mathbf{v}}(\mathbf{x})\|_{1}\leq\\ \sqrt{m}\sum_{i=1}^{n}\min\left\{\|\mathbf{C}_{:,i}\|\|(\mathbf{a}_{i}^{T},b_{i})\|,\frac{l_{k(i)}}{N_{min}}+\|\epsilon_{:,i}\|_{F}\right\}\\ +\sum_{j=1}^{m}\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
where:
  • $N_{j}$ is the set of null neurons with respect to output $j$.
  • $k(i)$ is the cluster of neuron $i$.
  • $N_{min}$ is the minimum cardinality of the non-null generators of a cluster.
  • $l_{k}$ is the objective value of the optimization criterion for cluster $k$.
  • $\epsilon_{j,i}$ is the difference/error between $c_{ji}(\mathbf{a}_{i}^{T},b_{i})$ and the cluster mean $\frac{\sum_{i\in I_{jk}}c_{ji}(\mathbf{a}_{i}^{T},b_{i})}{|I_{jk}|}$.
  • $\epsilon_{:,i}=[\epsilon_{1,i},\dots,\epsilon_{m,i}]$.
  • $\|\cdot\|_{F}$ is the Frobenius norm.
Heuristic Improvements
We obtain an advantage by dropping the bias from the clustering vectors – which leads to better upper hull approximations – and normalizing the vectors used for clustering. These two heuristics are discussed in detail in the appendix.
Deep and Convolutional Networks.
For deep networks, we can view each hidden layer of the network as an instance for our algorithm, having an input from the previous layer and providing an output to the next layer. Hence, we can recursively apply our multi-output algorithm to compress each hidden layer. For convolutional networks, we unravel the kernels of the weight tensors row-wise for $\mathbf{A}$ and column-wise for $\mathbf{C}$, apply our algorithm, and reshape them back to 4D tensors. If we also have batch normalization, then this can be dealt with by fusing the operations of batch normalization into the preceding convolutional or linear layer.
Non-uniform compression.
In Algorithm 2 we intentionally did not mention how clustering is performed. We can either use K-Means with a constant pruning ratio for all layers, or we can use hierarchical clustering with a global threshold parameter, like in the CUP framework. When used with hierarchical clustering, TropNNC differs from CUP in step 1, where we choose a different approach to build the clustering vectors (or filter features as Duggal et al. (2021) call them) of a convolutional layer, and more importantly in step 3, where we choose a different cluster representative based on tropical geometry. We also propose a modification to step 2 of CUP. Since the clustering vectors of different layers have varying dimensions, and vectors in higher-dimensional spaces tend to be more spread out, we introduce two variants for selecting the distance threshold for each layer:
  • Variant 1: For each layer, take the distance threshold to be some global constant times the square root of the dimension of the clustering vectors of the layer.
  • Variant 2: For each layer, take the distance threshold to be some global constant times the mean of the norms of the clustering vectors of the layer.
Limitations.
Limitations are discussed in the appendix.
5 Related Work
Our work falls under the category of structured pruning (He and Xiao, 2024). Here, approaches include data-free norm-based pruning (Li et al., 2017; He et al., 2019), KL-divergence based pruning (Luo and Wu, 2020), data-driven activation-based pruning (Lin et al., 2020; Sui et al., 2021), and data-driven filter pruning based on reconstruction loss minimization. Within the latter category, ThiNet (Luo et al., 2017) is a representative method that greedily removes filters contributing least to the next layer’s input, based on a data-driven reconstruction objective. Our approach resembles ThiNet in its structured layer-wise compression, but unlike ThiNet, it is data-free. Although layer-wise pruning is increasingly being replaced by global strategies – such as NISP (Yu et al., 2018) and SPvR (Hassan et al., 2025), which consider the importance of filters relative to the final layer – ThiNet remains a strong structured data-driven baseline to which we compare our method. More recently, data-free methods have been proposed. For example, CUP (Duggal et al., 2021) hierarchically clusters similar channels and selects representatives by $L^{1}$ norm. Our method builds on the same intuition of redundancy reduction via clustering, but introduces three key extensions: (i) modifying the clustering vectors, (ii) adapting the distance thresholds of hierarchical clustering to layer dimensions, and (iii) using tropical geometry to select representatives.
We built on Misiakos et al. (2022), who used a discrete form of the Hausdorff distance to geometrically bound the error between tropical polynomials, framing network reduction as a zonotope order reduction problem, and proposing a basic clustering-based compression method. Their method clusters similar neurons and replaces each cluster with the representative obtained by taking the mean of the input and output weights of the cluster. We refined their central theorem using the standard Hausdorff distance and introduce a novel algorithm that formulates compression as a simultaneous zonotope approximation problem. Our method makes a better initial estimation of the compressed network’s weights, and refines them by an iterative process. In the single-output case, our approach leads to stronger theoretical guarantees. Our method also has support for convolutional layers.
6 Experiments
Baselines.
We empirically evaluate TropNNC against baselines that perform structured pruning that do not require retraining. Specifically, as baselines we choose Neural Path K-Means by Misiakos et al. (2022), ThiNet, CUP, and the simple random and L1 structured pruning baselines. Neural Path K-Means is originally designed for linear layers only. To enable a fair comparison, we extended their approach using our proposed technique, making them applicable to convolutional layers as well. For the non-uniform variant of our framework, we compare it with CUP. As explained, our proposed algorithm enhances all three steps of CUP. In the presented experiments, we compare CUP exclusively with the fully enhanced version of TropNNC. To eliminate randomness, all random methods were performed 5 times on each model and the best performing compressed model was selected for testing. All metrics report the average over 5 repetitions of the same experiments with error bars indicating the standard deviation.
Datasets and networks.
We evaluate our framework on the MNIST (Deng, 2012) and CIFAR (Krizhevsky and Hinton, 2009) datasets, testing it across various models including simple multi-layer perceptrons (MLPs), convolutional neural networks (CNNs), AlexNet (Krizhevsky et al., 2012), and VGG (Simonyan and Zisserman, 2015). The non-uniform variant of our algorithm is tested on the CIFAR and ImageNet (Deng et al., 2009) datasets, with its performance evaluated across models such as VGG, and ResNet18 (He et al., 2016). The selected datasets and models are widely used benchmarks for neural network compression, and offer diversity in their complexity. We compare the methods based on test accuracies and FLOPs reduction, following the literature.
Experimental setup.
Only a single hyper-parameter needed tuning: ThiNet’s training samples used. For each dataset, we gradually increased this number until no further improvement of ThiNet was observed. $5000$ was found to suffice for both MNIST and CIFAR. The same number can be found by using the method presented in (Luo et al., 2017) for selecting the parameter.
6.1 MNIST and Fashion-MNIST Datasets
The first experiment is performed on the MNIST and Fashion-MNIST datasets. Table 1 compares TropNNC with Neural Path K-means for the same CNN network and the same pruning ratios as in (Misiakos et al., 2022). We compress the final linear layer. As shown in the results, for both datasets, our algorithm outperforms Neural Path K-means.
To evaluate the performance of our algorithm in compressing linear layers of deeper networks, we applied TropNNC to "deepNN", a MLP with layer sizes $28\times 28,512,256,128$, and $10$. The performance plots are provided in Figures 4(a) and 4(b). As illustrated, Neural Path K-Means achieves results comparable to L1-structured pruning. By contrast, TropNNC outperforms both Neural Path K-means and ThiNet.
To assess the performance of our algorithm in compressing convolutional layers, we applied TropNNC to "deepCNN2D", a LeNet-type convolutional neural network with ReLU activations. The performance plots are provided in Figures 4(c) and 4(d). The results demonstrate that TropNNC significantly outperforms Neural Path K-means, whose effectiveness appears reduced. Moreover, TropNNC matches, or even surpasses, ThiNet.
Table 1: Comparison of Neural Path K-means and TropNNC on MNIST and Fashion-MNIST
Percentage of Remaining Neurons MNIST Fashion-MNIST
Neural Path K-means TropNNC Neural Path K-means TropNNC
100.0 98.54 $\pm$ 0.16 98.54 $\pm$ 0.16 89.16 $\pm$ 0.21 89.16 $\pm$ 0.21
50.0 97.85 $\pm$ 0.39 98.49 $\pm$ 0.14 88.17 $\pm$ 0.46 89.00 $\pm$ 0.25
25.0 96.69 $\pm$ 1.06 98.36 $\pm$ 0.14 86.33 $\pm$ 0.87 88.70 $\pm$ 0.22
10.0 96.25 $\pm$ 1.39 97.96 $\pm$ 0.35 84.91 $\pm$ 1.28 88.24 $\pm$ 0.40
5.0 95.17 $\pm$ 2.36 97.06 $\pm$ 0.73 81.48 $\pm$ 3.90 87.42 $\pm$ 0.46
(a) deepNN, MNIST
(b) deepNN, F-MNIST
(c) deepCNN2D, MNIST
(d) deepCNN2D, F-MNIST
Figure 4: Compression of linear and convolutional layers of ReLU neural networks on MNIST datasets.
6.2 CIFAR-10 and -100 Datasets
In this experiment, we compress AlexNet and VGG trained on the CIFAR-10 and CIFAR-100 datasets to assess the performance of each compression method. Figures 5(a) and 5(b) illustrate the compression of the linear layers of AlexNet on CIFAR-10 and CIFAR-100, respectively. Additionally, Figures 5(c) and 5(d) show the compression of VGG’s convolutional layers for these datasets.
The results indicate that for larger datasets, TropNNC consistently outperforms Neural Path K-means, whose performance approaches that of random pruning. Furthermore, it matches or even surpasses the performance of ThiNet for the compression of linear layers. These findings highlight the effectiveness of TropNNC in handling more complex and larger-scale data scenarios. For the compression of convolutional layers of VGG, TropNNC matches ThiNet.
(a) AlexNet, lin., CIFAR10
(b) AlexNet, lin., CIFAR100
(c) VGG, conv., CIFAR10
(d) VGG, conv., CIFAR100
Figure 5: Compression of linear layers of AlexNet and convolutional layers of VGG on CIFAR datasets.
6.3 Non-uniform Pruning
We evaluate the effectiveness of the non-uniform variant of TropNNC. We compress various models, such as VGG on CIFAR-10, and ResNet18 on ImageNet. The results are summarized in Tables 2 and 3. We should emphasize that we did not apply any form of fine-tuning or re-training. Table 4 provides an ablation of the iterative variant of TropNNC.
Our findings indicate that our method demonstrates a clear advantage. Our approach shows a significant performance improvement for the VGG model, while the benefits are comparatively modest for ResNet18. The substantial advantage observed with VGG remains somewhat unexplained, and we have yet to determine why this effect is less pronounced for the ResNet architecture. Further investigation may be required to fully understand these discrepancies and to optimize our approach across different model types.
We also found that the first variant of non-uniform TropNNC excelled at reducing the overall network size but was less effective at minimizing inference operations. In contrast, the second variant performed well in both tasks, outperforming CUP. Upon further analysis, we concluded that the first variant tends to focus more aggressively on the final layers, where parameter count is high due to the increased number of channels, but the number of operations is lower because of smaller image sizes. Meanwhile, the second variant, like CUP, also targets the initial layers, where fewer parameters are present, but a greater number of operations is required.
Table 2: Comparison of CUP and TropNNC (variant 1) accuracy across different pruning thresholds on CIFAR10, VGG.
Method Threshold #params $\downarrow$ FLOPS $\downarrow$ Acc. $\uparrow$
Original - 14.7M 0.63G 93.64
CUP 0.15 6.24M 0.46G 92.41
TropNNC (3 iters) 0.014 5.22M 0.46G 93.58
CUP 0.20 4.62M 0.42G 67.91
TropNNC (3 iters) 0.017 3.61M 0.42G 93.02
CUP 0.25 3.61M 0.39G 16.02
TropNNC (3 iters) 0.02 2.75M 0.39G 91.71
Table 3: Comparison of CUP and TropNNC accuracy across different pruning thresholds on ImageNet, ResNet18.
Method Threshold #params $\downarrow$ FLOPS $\downarrow$ Acc. $\uparrow$
Original - 11.69M 3.64G 69.75
CUP 0.5 11.66M 3.58G 66.80
TropNNC (v2) 1.1 11.66M 3.48G 68.95
CUP 0.6 11.49M 3.38G 53.40
TropNNC (v2) 1.2 11.46M 3.25G 61.30
CUP 0.65 11.17M 3.15G 28.45
TropNNC (v2) 1.25 10.98M 2.92G 41.65
TropNNC (v1) 0.0187 9.84M 3.46G 59.15
Table 4: Comparison of non-uniform TropNNC and iterative TropNNC (variant 1) accuracy on CIFAR10, VGG.
Method Threshold #params $\downarrow$ FLOPS $\downarrow$ Acc. $\uparrow$
Original - 14.7M 0.63G 93.64
TropNNC 0.02 2.75M 0.39G 86.2
TropNNC (3 iters) 0.021 2.52M 0.38G 91.2
TropNNC 0.025 1.80M 0.33G 34.89
TropNNC (3 iters) 0.0255 1.76M 0.33G 69.95
7 Conclusion
We proposed TropNNC, a tropical geometrical method for structured, data-free pruning of linear and convolutional layers of ReLU activated neural networks. Based on clustering similar neurons/channels, it selects better cluster representatives than previous work by using tropical geometry and optimization, with provable compression guarantees. TropNNC significantly outperforms prior work on tropical geometric pruning and manages to match or even surpass the performance of the data-driven ThiNet. In non-uniform pruning it outperforms CUP, with significant improvement in the case of the VGG architecture. Our findings highlight the potential of tropical geometry in the realm of neural network compression.
Acknowledgments
The research project was supported by the Hellenic Foundation for Research and Innovation (H.F.R.I.) under the “2nd Call for H.F.R.I. Research Projects to support Faculty Members & Researchers” (Project Number:2656, Acronym: TROGEMAL).
Appendix
In this appendix, we first go into more detail regarding the heuristic improvements of our algorithm. Afterwards, we introduce the background of tropical algebra which was omitted from the main text. Finally, we provide the proofs for the theorems and propositions stated in the main text.
Before we proceed with the technical content of this appendix, we first provide an ethical statement for this work and a disclosure of the limitations of our method.
Ethical statement
This work does not introduce any new ethical concerns beyond those already present in prior research on network pruning and model compression. As with existing pruning methods, our approach aims to reduce model size and computational cost, and does not involve the use of sensitive data, human subjects, or deployment in high-risk applications.
Limitations
A key limitation of our method is that it relies on the presence of consecutive convolutional or linear layers – that is, the existence of intermediate hidden layers. In addition, it only holds for networks with ReLU activations.
Heuristic Improvements
In the experiments comparing TropNNC and CUP, TropNNC was enhanced with 2 heuristic improvements. These improvements are not rigorously backed by theory; however, we try to provide intuitive explanations as to why they work.
Heuristic 1.
First, let us focus on the single-output case, where the problem of network compression reduces to the problem of zonotope approximation. In this setting, Algorithm 1 suggests clustering similar generators based on their $L^{2}$ distances and replacing each cluster with the sum of its elements. The intuition is as follows: if the generators within a cluster are exactly equal, then through their Minkowski sum, these generators become equivalent to a single generator given by their sum. We can take this one step further: if we have generators that are parallel to each other, their Minkowksi sum reduces to a single equivalent generator without introducing any error. This suggest that generator similarity is determined by direction (i.e. cosine similarity) and not $L^{2}$ distance. Hence, an improvement to the single-output algorithm is to cluster the generators based on cosine similarity. In practice, this can be achieved by applying K-Means clustering to normalized generators, while still using the original generators in the summation step.
A similar heuristic can also be applied to the multi-output algorithm. By normalizing the input weights of the clustering vectors (i.e. the vectors used for clustering similar neurons/channels, not the weights used for calculating the representative of each cluster) we obtain an improvement to the algorithm.
Heuristic 2.
The second heuristic is motivated by Theorem 9, which states that two tropical polynomials are functionally identical if the upper hulls of their extended Newton polytopes coincide (i.e. only the vertices of the upper hull contribute to non-redundant terms). Indeed, this theorem also serves as motivation for Theorem 2, which shows that approximate equality of the extended Newton polytopes implies approximate functional equivalence of the corresponding tropical polynomials. Although we do not formally prove it, this theorem can be further refined by bounding the error of two tropical polynomials by the Hausdorff distance of the upper hulls of their extended Newton polytopes:
$)\frac{1}{\rho}\max_{\mathbf{x}\in B}|p(\mathbf{x})-\tilde{p}(\mathbf{x})|\leq H(UF(P),UF(\tilde{P}$))
Based on this refined result, it becomes clear that to compress networks, it suffices to approximate the upper hulls of the zonotopes of the network, rather than the zonotopes themselves.
We can take this idea one step further. Consider the upper hull of an extended Newton polytope, and project it by eliminating the bias term. Then, the points on the convex hull of the resulting projection (which coincides with the convex hull of the (non-extended) Newton polytope) correspond to the slopes of the outer linear regions of the polynomials, i.e. the regions which extend to infinity. We can view these points as being the "outer" points of the upper hull. When approximating a tropical polynomial, these unbounded regions are particularly important, because they have the potential to make the sup-norm of the approximation error infinite (something which is not captured by Theorem 2 because we divide by $\rho$ and take the $L^{1}$-norm error on a bounded set).
Based on this insight, our goal is to design an algorithm that prioritizes the approximation of the upper hulls of the zonotopes – particularly the vertices corresponding to the unbounded regions of the tropical polynomials. (i.e. the "outer" vertices of the upper hulls).
Let us first see how vanilla TropNNC behaves on a simple example. Consider the tropical polynomial $v(x)=\max(0,-x+5)+\max(0,x+5)+\max(0,x)$. Such a polynomial and its corresponding zonotope are depicted in Figure 6(a). Notice that the upper hull has 4 points, each corresponding to a linear region of the polynomial. It also has 2 "outer" points, which correspond to the 2 outer regions of the polynomial extending to infinity.
We reduce the terms from 3 to $K=2$. First, TropNNC performs clustering on the generators of the zonotope, forming the clusters $\{(-1,5),(1,5)\}$ and $\{(1,0)\}$. It then replaces the first cluster with the representative $(0,10)$ and the second cluster with the representative $(1,0)$. The reduced tropical polynomial is $\tilde{v}(x)=\max(0,10)+\max(0,x)=10+\max(0,x)$. The compression procedure and the corresponding reduced function is depicted in Figure 6(b).
It is evident that this approximation is not ideal: while the reduced polynomial accurately captures the two central regions, it fails to preserve the unbounded regions. This causes the approximation error becomes unbounded.
Since we have $K=2$, the best possible approximation using only 2 terms is $v^{\prime}(x)=\max(0,-x+5)+\max(0,2x+5)$. This approximation captures both the important outer regions with zero error, and also one of the central regions. The question however remains: how could we have arrived at such an approximation? To answer this, notice that $v^{\prime}(x)=\max(0,-x+5)+\max(0,(x+5)+(x))$. Thus, if TropNNC had initially formed the clusters $\{(-1,5)\}$ and $\{(1,5),(1,0)\}$, then it would have been able to find the better approximation. Now, the modification we have to make becomes clear: instead of forming the clusters based on the generators themselves, we can form the clusters based on the reduced generators with the bias suppressed (and, of course, for the summation step the bias remains). The modified compression procedure is depicted in Figure 6(c). The new approximation clearly provides a closer fit to the upper hull and captures the “outer” points exactly, resulting in zero error in the outer unbounded regions.
To conclude, another heuristic improvement comes from dropping the bias from the clustering vectors, seeing as this helps the algorithm achieve better upper hull approximations and prioritize the important "outer" vertices of the upper hull.
(a) Example tropical polynomial $v(x)=\max(0,-x+5)+\max(0,x+5)+\max(0,x)$ and corresponding zonotope.
(b) Approximation using vanilla TropNNC.
(c) Approximation using Heuristic 2.
Figure 6: Example of reduction of tropical polynomial.
Background on Tropical Algebra
We begin with the definitions of tropical polynomials, tropical rational functions, Newton polytopes, and zonotopes.
Tropical Polynomials and Rational Functions.
Within the max-plus semiring, we can define polynomials. A tropical polynomial $f$ in $d$ variables $\mathbf{x}=(x_{1},\ldots,x_{d})$ is defined as the function:
$f(\mathbf{x})=\bigvee_{i=1}^{n}\left\{\mathbf{a}_{i}^{T}\mathbf{x}+b_{i}\right\}=\max_{i\in[n]}\left\{\mathbf{a}_{i}^{T}\mathbf{x}+b_{i}\right\},$ (4)
where $n]:=\{1,...,n\}$. Here, $n$ represents the rank of the tropical polynomial. Each monomial term $\{\mathbf{a}_{i}^{T}\mathbf{x}+b_{i}\}$ of the polynomial has an exponent or slope $\mathbf{a}_{i}\in\mathbb{R}^{d}$ and a coefficient or bias $b_{i}\in\mathbb{R}$. Each monomial term corresponds to a plane in $\mathbb{R}^{d+1}$. Consequently, tropical polynomials are piecewise linear convex functions. Specifically, every tropical polynomial is a continuous piecewise linear convex function, and every continuous piecewise linear convex function can be expressed (though not uniquely) as a tropical polynomial (Maclagan and Sturmfels, 2021). The set of tropical polynomials in $\mathbf{x}$ defines the semiring $\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$. Figures 7(a), 7(b) illustrate examples of tropical polynomials in one and multiple variables, respectively.
Tropical rational functions are defined as the difference of two tropical polynomials $p$ and $q$:
$r(\mathbf{x})=p(\mathbf{x})-q(\mathbf{x})$
Tropical rational functions correspond to general piecewise linear functions. Specifically, every tropical rational function is a continuous piecewise linear function, and every continuous piecewise linear function can be expressed (though not uniquely) as a tropical rational function. Figures 7(c), 7(d) provide examples of tropical rational functions.
Figure 7: (a) depicts a single-variate tropical polynomial, (b) depicts a multi-variate tropical polynomial, (c) depicts a single-variate tropical rational function, (d) depicts a multi-variate tropical rational function
Newton Polytopes.
For a tropical polynomial $f$ as defined in (4), we define its Newton polytope as the convex hull of the slopes $\mathbf{a}_{i}$ of $f$
$\mathrm{Newt}(f):=\mathrm{conv}\left\{\mathbf{a}_{i},i\in[n]\right\}.$
Additionally, we define the extended Newton polytope of a tropical polynomial $f$ as the convex hull of the slopes $\mathbf{a}_{i}$ of $f$ extended in their last dimension by the coefficient $b_{i}$
$\mathrm{ENewt}(f):=\mathrm{conv}\left\{(\mathbf{a}_{i}^{T},b_{i}),i\in[n]\right\}.$
The following proposition allows us to calculate the (extended) Newton polytope of expressions of tropical polynomials, and it was used in the main text to deduce that the extended Newton polytopes of the outputs of the network are zonotopes.
Proposition 7 (Zhang et al., 2018).
Let $f,g\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$ be two tropical polynomials in $\mathbf{x}$. For the extended Newton polytope, the following holds:
$\mathrm{ENewt}\left(f\vee g\right)$ $=\mathrm{conv}\left\{\mathrm{ENewt}(f)\cup\mathrm{ENewt}(g)\right\}$
$\mathrm{ENewt}(f+g)$ $=\mathrm{ENewt}(f)\oplus\mathrm{ENewt}(g).$
The Minkowski sum $\oplus$ of two polytopes (or more generally subsets of $\mathbb{R}^{d}$) $P$ and $Q$ is defined as:
$P\oplus Q:=\{p+q\mid p\in P,q\in Q\}.$
Using the principle of induction, Proposition 7 can be generalized to any finite tropical expression of tropical polynomials.
The upper envelope or upper hull $UF(P)$ of an extended Newton polytope is defined as the set of all points $(\mathbf{a}^{T},b)$ of the polytope $P$ that are not "shadowed" by any other part of the polytope when viewed from above (last dimension). This means that there is no $b^{\prime}>b$ such that $(\mathbf{a}^{T},b^{\prime})$ belongs to $P$. We have the following useful lemma.
Lemma 8.
Let $p\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$ a tropical polynomial in $d$ variables with extended Newton polytope $P=\mathrm{ENewt}(p)$. If $(\mathbf{a}^{T},b)$ lies below the upper envelope of $P$, then $\forall\mathbf{x}\in\mathbb{R}^{d},\mathbf{a}^{T}\mathbf{x}+b\leq p(\mathbf{x})$. The inequality is strict if $(\mathbf{a}^{T},b)$ lies strictly below the upper envelope.
Proof.
Since $(\mathbf{a}^{T},b)$ lies below the upper envelope of $P$, there exists a point
$(\mathbf{a}^{T},b^{\prime})=\sum_{i=1}^{k}\lambda_{i}\mathbf{v}_{i},\quad\sum_{i=1}^{k}\lambda_{i}=1$
on a face of the upper envelope of $P$ defined by the points $\mathbf{v}_{1},\ldots,\mathbf{v}_{k}\in V_{UF(P)}$ such that $b^{\prime}\geq b$. Therefore, we have:
$\mathbf{a}^{T}\mathbf{x}+b\leq\mathbf{a}^{T}\mathbf{x}+b^{\prime}=\langle(\mathbf{a}^{T},b^{\prime}),(\mathbf{x},1)\rangle$
$\sum_{i=1}^{k}\lambda_{i}\langle\mathbf{v}_{i},(\mathbf{x},1)\rangle\leq\max_{i=1,\ldots,k}\langle\mathbf{v}_{i},(\mathbf{x},1)\rangle\leq p(\mathbf{x}).$
If $(\mathbf{a}_{i}^{T},b_{i})$ lies strictly below the upper envelope of $P$, then $b^{\prime}>b$, and the inequality is strict. ∎
The extended Newton polytope provides a geometrical interpretation for studying tropical polynomials. For instance, the following theorem holds.
Theorem 9.
For any two tropical polynomials $f,g\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$, the following holds:
$f=g\Leftrightarrow UF(\mathrm{ENewt}(f))=UF(\mathrm{ENewt}(g))$
This implies that two tropical polynomials are functionally identical if and only if their extended Newton polytopes have the same upper envelope.
The above theorems indicate that a tropical polynomial is fully functionally determined by the upper envelope of its extended Newton polytope, as shown by the following example.
Example 2.
Consider the polynomials corresponding to Figure 8.
$f(x,y)=\max\{0,-y+1,y+1\}$
$g(x,y)=\max\{x+1,-x+1\}.$
We have that
$(f\vee g)(x,y)=\max\{0,x+1,y+1,-y+1,-x+1\}$
$(f+g)(x,y)=\max\{x+1,-x+1,x-y+2,$
xy+2,x+y+2,x+y+2}.-x-y+2,x+y+2,-x+y+2\}.
The extended Newton polytopes of $f,g,f\vee g,f+g$ are shown in Figure 8. The polynomial $f\vee g$ can be reduced as follows:
$(f\vee g)(x,y)=\max\{x+1,y+1,-y+1,-x+1\},$
which corresponds to the vertices of the upper envelope of $\mathrm{ENewt}(f\vee g)$.
(a) $f$
(b) $g$
(c) $f\vee g$
(d) $f+g$
Figure 8: Operations on tropical polynomials. $\mathrm{ENewt}(f\vee g)$ corresponds to the convex hull of the union of the vertices of the polytopes $\mathrm{ENewt}(f),\mathrm{ENewt}(g)$. $\mathrm{ENewt}(f+g)$ corresponds to the Minkowski sum of $\mathrm{ENewt}(f),\mathrm{ENewt}(g)$. For the polytope $\mathrm{ENewt}(f\vee g)$ we illustrate with blue the upper envelope, which consists of a single face. The vertices of the upper envelope are the only non reduntant terms of the polynomial $f\vee g$.
Zonotopes.
In polytope theory, zonotopes are a special class of convex polytopes that can be defined as the Minkowski sum of a finite set of line segments (or edges). Formally, given a set of line segments $g_{1},\ldots,g_{n}$, the zonotope is defined as
$Z:=\bigoplus_{i\in[n]}g_{i}$
The line segments $g_{i}$ are referred to as the zonotope’s generators.
Alternatively, a zonotope can be expressed equivalently by a set of vectors $\mathbf{v}_{1},\ldots,\mathbf{v}_{n}\in\mathbb{R}^{d}$ and a starting point $\mathbf{s}\in\mathbb{R}^{d}$. By taking the generators to be the segments $\mathbf{0},\mathbf{v}_{1}],\ldots,[\mathbf{0},\mathbf{v}_{n}]$ and translating the first segment by $\mathbf{s}$, we obtain the equivalent form:
$Z=\left\{\mathbf{s}+\sum_{i=1}^{n}\lambda_{i}\mathbf{v}_{i}\mid 0\leq\lambda_{i}\leq 1\right\}$
In this context, the vectors $\mathbf{v}_{i}$ are sometimes referred to as the zonotope’s generators, meaning the segments $\mathbf{0},\mathbf{v}_{i}]$. When the starting point $\mathbf{s}$ is not mentioned, it is assumed to be the origin $\mathbf{0}$.
For a zonotope with a starting point $\mathbf{s}\in\mathbb{R}^{d}$ and generators $\mathbf{v}_{1},\ldots,\mathbf{v}_{n}\in\mathbb{R}^{d}$, a vertex $\mathbf{u}\in V_{Z}$ corresponds to points where $\lambda_{i}=0\ \text{or}\ 1$. The vertex $\mathbf{u}$ can be expressed as:
$\mathbf{u}=\mathbf{s}+\sum_{i\in I}\mathbf{v}_{i},$
where $I\subseteq[n]$.
Tropical polynomial approximation based on Hausdorff distance
Before we proceed with the proof of our refined Theorem 2, we first provide an auxiliary lemma, stated implicitly in the main text, which is necessary to see how our bound is indeed tighter than the bound of Misiakos et al. (2022). This lemma is also used for the proofs of Theorem 2, and Propositions 4 and 6.
Lemma 10.
Due to the convexity and compactness of polytopes, we have that
$H(P,\tilde{P})=\max\left\{\max_{\mathbf{u}\in V_{P}}\mathrm{dist}(\mathbf{u},\tilde{P}),\max_{\mathbf{v}\in V_{\tilde{P}}}\mathrm{dist}(P,\mathbf{v})\right\}$
Proof.
We will prove that
[…]
i.e., the supremum is attained at some vertex of $P$.
The polytope $P$, the domain of the supremum, is convex and compact. Thus, it suffices to prove that the function
$f(\mathbf{u})=\operatorname{dist}(\mathbf{u},\tilde{P})=\inf_{\tilde{\mathbf{u}}\in\tilde{P}}\|\mathbf{u}-\tilde{\mathbf{u}}\|$
is convex in terms of $\mathbf{u}$.
Let $\mathbf{u}_{1},\mathbf{u}_{2}\in P$. By the compactness of $\tilde{P}$, there exist points $\tilde{\mathbf{u}}_{1},\tilde{\mathbf{u}}_{2}\in\tilde{P}$ such that $f(\mathbf{u}_{1})=\operatorname{dist}(\mathbf{u}_{1},\tilde{P})=\|\mathbf{u}_{1}-\tilde{\mathbf{u}}_{1}\|$ and $f(\mathbf{u}_{2})=\operatorname{dist}(\mathbf{u}_{2},\tilde{P})=\|\mathbf{u}_{2}-\tilde{\mathbf{u}}_{2}\|$.
For every $\lambda\in[0,1]$, we have that
$\lambda f(\mathbf{u}_{1})+(1-\lambda)f(\mathbf{u}_{2})$ $=\lambda\|\mathbf{u}_{1}-\tilde{\mathbf{u}}_{1}\|+(1-\lambda)\|\mathbf{u}_{2}-\tilde{\mathbf{u}}_{2}\|$
$\geq\|\lambda(\mathbf{u}_{1}-\tilde{\mathbf{u}}_{1})+(1-\lambda)(\mathbf{u}_{2}-\tilde{\mathbf{u}}_{2})\|$
$=\|\lambda\mathbf{u}_{1}+(1-\lambda)\mathbf{u}_{2}-\lambda\tilde{\mathbf{u}}_{1}-(1-\lambda)\tilde{\mathbf{u}}_{2}\|.$
By the convexity of $\tilde{P}$, $\tilde{\mathbf{u}}=\lambda\tilde{\mathbf{u}}_{1}+(1-\lambda)\tilde{\mathbf{u}}_{2}\in\tilde{P}$. Hence,
$\lambda f(\mathbf{u}_{1})+(1-\lambda)f(\mathbf{u}_{2})\geq\|\lambda\mathbf{u}_{1}+(1-\lambda)\mathbf{u}_{2}-\tilde{\mathbf{u}}\|\geq f(\lambda\mathbf{u}_{1}+(1-\lambda)\mathbf{u}_{2}),$
which concludes the proof. ∎
We continue with the proof of Theorem 2.
Proof (Theorem 2).
Consider a point $\mathbf{x}\in B$ and assume that $p(\mathbf{x})=\mathbf{a}^{T}\mathbf{x}+b$ and $\tilde{p}(\mathbf{x})=\mathbf{c}^{T}\mathbf{x}+d$. Take an arbitrary $(\mathbf{u}^{T},v)\in\tilde{P}$. This point lies below the upper envelope of $\tilde{P}$. Thus, by Lemma 8, we have that $\tilde{p}(\mathbf{x})\geq\mathbf{u}^{T}\mathbf{x}+v$. Choose $(\mathbf{u}^{T},v)$ to be the closest point to $(\mathbf{a}^{T},b)$. Then,
$p(\mathbf{x})-\tilde{p}(\mathbf{x})$ $\leq p(\mathbf{x})-(\mathbf{u}^{T},v)\begin{pmatrix}\mathbf{x}\\ 1\end{pmatrix}$
$=((\mathbf{a}^{T},b)-(\mathbf{u}^{T},v))\begin{pmatrix}\mathbf{x}\\ 1\end{pmatrix}$
$\leq\left\|(\mathbf{a}^{T},b)-(\mathbf{u}^{T},v)\right\|\left\|\begin{pmatrix}\mathbf{x}\\ 1\end{pmatrix}\right\|$
$\leq\operatorname{dist}((\mathbf{a}^{T},b),\tilde{P})\cdot\rho$
$\leq\max_{(\mathbf{a}^{T},b)\in V_{P}}\operatorname{dist}((\mathbf{a}^{T},b),\tilde{P})\cdot\rho,$
where the second inequality is due to the Cauchy-Schwarz inequality.
In a similar manner, take an arbitrary $(\mathbf{r}^{T},s)\in P$. This point lies below the upper envelope of $P$. Thus, by Lemma 8, we have that $p(\mathbf{x})\geq\mathbf{r}^{T}\mathbf{x}+s$. Choose $(\mathbf{r}^{T},s)$ to be the closest point to $(\mathbf{c}^{T},d)$. Then,
$p(\mathbf{x})-\tilde{p}(\mathbf{x})$ $\geq(\mathbf{r}^{T},s)\begin{pmatrix}\mathbf{x}\\ 1\end{pmatrix}-\tilde{p}(\mathbf{x})$
$=((\mathbf{r}^{T},s)-(\mathbf{c}^{T},d))\begin{pmatrix}\mathbf{x}\\ 1\end{pmatrix}$
$\geq-\left\|(\mathbf{r}^{T},s)-(\mathbf{c}^{T},d)\right\|\left\|\begin{pmatrix}\mathbf{x}\\ 1\end{pmatrix}\right\|$
$\geq-\operatorname{dist}((\mathbf{r}^{T},s),\tilde{P})\cdot\rho$
$\geq-\max_{(\mathbf{c}^{T},d)\in V_{\tilde{P}}}\operatorname{dist}((\mathbf{c}^{T},d),P)\cdot\rho.$
Finally, we obtain that
$\max_{(\mathbf{c}^{T},d)\in V_{\tilde{P}}}\operatorname{dist}((\mathbf{c}^{T},d),P)\cdot\rho\leq p(\mathbf{x})-\tilde{p}(\mathbf{x})\leq\max_{(\mathbf{a}^{T},b)\in V_{P}}\operatorname{dist}((\mathbf{a}^{T},b),\tilde{P})\cdot\rho,$
which implies
$\frac{1}{\rho}\lvert p(\mathbf{x})-\tilde{p}(\mathbf{x})\rvert\leq\max\left\{\max_{(\mathbf{a}^{T},b)\in V_{P}}\operatorname{dist}((\mathbf{a}^{T},b),\tilde{P}),\max_{(\mathbf{c}^{T},d)\in V_{\tilde{P}}}\operatorname{dist}((\mathbf{c}^{T},d),P)\right\},\quad\forall\mathbf{x}\in B.$
Therefore, by Lemma 10, we have,
[…]
Proofs of Proposition 4 and Corollary 5
We continue with the proofs of Proposition 4 and Corollary 5.
Proof (Prop. 4).
For the output function, it holds that
$v(\mathbf{x})=p(\mathbf{x})-q(\mathbf{x}),\quad\tilde{v}(\mathbf{x})=\tilde{p}(\mathbf{x})-\tilde{q}(\mathbf{x}).$
From the triangle inequality, we deduce
[…]
Thus, it suffices to get a bound on H(P,P~)H($\tilde{P}$) and H(Q,Q~)H($\tilde{Q}$).
Let $I_{+}\subseteq[n]$ and $I_{-}\subseteq[n]$ be the sets of positive and negative generators, respectively. First, we deal with H(P,P~)H($\tilde{P}$). Notice that $\forall\mathbf{v}\in V_{\tilde{P}},\mathbf{v}$ is the sum of generators of some clusters of $P$. Thus $\forall\mathbf{v}\in V_{\tilde{P}},\mathbf{v}$ is a vertex of $P$. Hence,
$\operatorname{dist}(P,\mathbf{v})=0,\quad\forall\mathbf{v}\in V_{\tilde{P}}$.
Let $I_{k}\subseteq[n]$ be the set of generators that belong to cluster $k$. Let $C_{+}\subseteq[K]$ be the set of clusters of positive generators, and $C_{-}\subseteq[K]$ the set of clusters of negative generators.
Consider any vertex $\mathbf{u}$ of $P$. This vertex can be written as the sum of generators $c_{i}(\mathbf{a}_{i}^{T},b_{i})$, for some subset $I_{+}^{\prime}\subseteq I_{+}$. Thus,
$\mathbf{u}=\sum_{i\in I_{+}^{\prime}}c_{i}(\mathbf{a}_{i}^{T},b_{i}).$
For every positive generator $c_{i}(\mathbf{a}_{i}^{T},b_{i})$ that belongs to cluster $k$, define
$x_{i}=\arg\min_{x}\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})-x\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|,$
i.e., project the generator onto its cluster representative. Since every inner-cluster pair of generators forms an acute angle, any generator and its cluster representative (sum of generators of the cluster) will also form an acute angle, and thus $x_{i}\geq 0$.
For every cluster $k\in C_{+}$ define $I_{k}^{\prime}=I_{k}\cap I_{+}^{\prime}$ and
$\tilde{x}_{k}=\sum_{i\in I_{k}^{\prime}}x_{i}.$
Since the cluster representative is the sum of the generators of the cluster, we have
$\sum_{i\in I_{k}}x_{i}=1\Rightarrow\tilde{x}_{k}=\sum_{i\in I_{k}^{\prime}}x_{i}\leq 1$
Thus, for every cluster $k\in C_{+}$, the point $\tilde{x}_{k}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ lies inside the segment $\mathbf{0},\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})]$ and thus belongs to $\tilde{P}$.
For the vertex $\mathbf{u}$, we choose to compare it with the point
$\sum_{k\in C_{+}}\tilde{x}_{k}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\in\tilde{P}$.
Thus, we have that
[…]
$\leq\sum_{k\in C_{+}}\left\|\sum_{i\in I_{k}^{\prime}}c_{i}(\mathbf{a}_{i}^{T},b_{i})-\tilde{x}_{k}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|$
$\leq\sum_{k\in C_{+}}\left\|\sum_{i\in I_{k}^{\prime}}\left[c_{i}(\mathbf{a}_{i}^{T},b_{i})-x_{i}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right]\right\|$
$\leq\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})-x_{i}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|$
$=\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\min_{x}\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})-x\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|$
$=\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\min_{x}\Bigl\|c_{i}(\mathbf{a}_{i}^{T},b_{i})-x|I_{k}|(c_{i}(\mathbf{a}_{i}^{T},b_{i})+\epsilon_{i})\Bigr\|$
$\leq\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\min_{x}\left\{|1-x|I_{k}||\cdot\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|+|x|I_{k}||\cdot\left\|\epsilon_{i}\right\|\right\}$
$=\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\left\|\epsilon_{i}\right\|\right\}$
$\leq\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}$
$=\sum_{i\in I_{+}^{\prime}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\},$
where $\epsilon_{i}$ is the error between generator $i$ and the cluster center/mean of K-means.
The maximum value of the upper bound occurs when $I_{+}^{\prime}=I_{+}$. Thus,
$\max_{\mathbf{u}\in V_{P}}\operatorname{dist}(\mathbf{u},\tilde{P})\leq\sum_{i\in I_{+}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}.$
Finally, we have
[…]
$\leq\max\left\{\sum_{i\in I_{+}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\},0\right\}$
$=\sum_{i\in I_{+}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}.$
Similarly, for H(Q,Q~)H($\tilde{Q}$), we have
$H(Q,\tilde{Q})\leq\sum_{i\in I_{-}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}.$
Combining, we get
$\frac{1}{\rho}\left|v(\mathbf{x})-\tilde{v}(\mathbf{x})\right|\leq\sum_{i\in I}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\},$
which concludes the proof of Proposition 4. ∎
Proof (Corollary 5).
The bound of Misiakos et al. (2022) is the following:
$\frac{1}{\rho}\max_{\mathbf{x}\in B}|v(\mathbf{x})-\tilde{v}(\mathbf{x})|\leq K\delta_{\max}+\left(1-\frac{1}{N_{\max}}\right)\sum_{i=1}^{n}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|.$
We will show that
$K\delta_{\max}+\left(1-\frac{1}{N_{\max}}\right)\sum_{i=1}^{n}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|\geq\sum_{i\in I}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}.$
This can be rewritten as
$K\delta_{\max}+\left(1-\frac{1}{N_{\max}}\right)\sum_{i=1}^{n}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|\geq\sum_{i\in I}\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|+\sum_{i\in I}\min\left\{0,\delta_{\max}-\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|\right\}.$
Further simplifying, we get
$K\delta_{\max}\geq\frac{1}{N_{\max}}\sum_{i=1}^{n}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|+\sum_{i=1}^{n}\min\left\{0,\delta_{\max}-\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|\right\}.$
It suffices to show that for every cluster $k$, we have:
$\delta_{\max}\geq\frac{1}{|I_{k}|}\sum_{i\in I_{k}}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|+\sum_{i\in I_{k}}\min\left\{0,\delta_{\max}-|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|\right\}.$
However, it holds that
$\sum_{i\in I_{k}}\min\left\{0,\delta_{\max}-|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|\right\}\leq\delta_{\max}-\max_{i\in I_{k}}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|,$
and
$\max_{i\in I_{k}}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|\geq\frac{1}{|I_{k}|}\sum_{i\in I_{k}}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|.$
Hence, we have
$\sum_{i\in I_{k}}\min\left\{0,\delta_{\max}-|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|\right\}+\frac{1}{|I_{k}|}\sum_{i\in I_{k}}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|\\ \leq\delta_{\max}-\max_{i\in I_{k}}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|+\frac{1}{|I_{k}|}\sum_{i\in I_{k}}|c_{i}|\|(\mathbf{a}_{i}^{T},b_{i})\|\leq\delta_{\max},$
which concludes the proof. ∎
Proof of Proposition 6
Before we proceed with the proof of Proposition 6, we first give the definition of null neurons and generators.
Definition 11 (Null neuron/generator).
A neuron/generator $i\in[n]$ that belongs to cluster $k\in[K]$ is a null neuron/generator with respect to output $j\in[m]$ if $\tilde{c}_{jk}c_{ji}\leq 0$. $N_{j}$ is the set of all null neurons with respect to output $j$.
Proof (Prop. 6).
Assume the algorithm iterative scheme has reached a stationary point (otherwise, assume the last step is an output weight update and the proof works fine). First, we focus on a single output, say $j$-th output. We will bound $H(P_{j},\tilde{P}_{j}),H(Q_{j},\tilde{Q}_{j})$ for all $j\in[m]$.
Let $I_{j+},I_{j-}$ be the sets of positive and negative generators of output $j$. Let $I_{k}$ be the set of neurons that belong to cluster $k$. Let $C_{j+}$ be the set of positive clusters for output $j$ (i.e. clusters for which $\tilde{c}_{jk}>0$), and $C_{j-}$ be the set of negative clusters for output $j$. Let $I_{jk}=I_{k}\cap I_{j+}$ if $k$ is a positive cluster, else $I_{jk}=I_{k}\cap I_{j-}$
Consider any vertex $\mathbf{u}$ of $P_{j}$. This vertex can be written as the sum of generators $c_{ji}(\mathbf{a}_{i}^{T},b_{i})$, for some subset $I_{j+}^{\prime}\subseteq I_{j+}$. Thus,
$\mathbf{u}=\sum_{i\in I_{j+}^{\prime}}c_{ji}(\mathbf{a}_{i}^{T},b_{i}).$
For every generator $c_{ji}(\mathbf{a}_{i}^{T},b_{i}),i\in I_{j+}$ that belongs to positive cluster $k\in C_{j+}$, define
$x_{ji}=\arg\min_{x}\left\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})-x\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|,$
i.e., project the generator onto $\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ of its cluster $k$.
Assume a variant of the algorithm, where the optimization criterion is the following:
$\sum_{j=1}^{m}\left\||\tilde{C}_{jk}|(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})-\sum_{i\in I_{jk}}|C_{ji}|(\mathbf{a}_{i}^{T},b_{i})\right\|^{2}.$
The sign of $\tilde{C}_{jk}$ never changes and gets fixed based on the initial solution. The set $I_{jk}$ of the non-null generators of cluster $k$ in terms of output $j$ depends on the sign of $\tilde{C}_{jk}$, and it is determined by the initial solution. The output weight update rule changes: We update the absolute value of the weight $|\tilde{C}_{jk}|$. The update is performed as normal if the result is positive, otherwise we set $|\tilde{C}_{jk}|=0$. We deduce that throughout the execution of the algorithm, after every output weight update step the following holds:
  • If $|\tilde{C}_{jk}|=0$ then, every generator of cluster $k$ is a null generator by definition.
  • If $|\tilde{C}_{jk}|>0$ then the following argument holds.
By hypothesis, for every cluster $k$, the vectors of the set $\{(\mathbf{a}_{i}^{T},b_{i})|i\in I_{k}\}$ form pair-wise acute angles. Every vector of the set of vectors $\{\sum_{i\in I_{jk}}|c_{ji}|(\mathbf{a}_{i}^{T},b_{i})|j\in[m]\}$ lies inside the cone of the set of vectors $\{(\mathbf{a}_{i}^{T},b_{i})|i\in I_{k}\}$. It is easy to verify that the representative $|\tilde{c}_{jk}|(\tilde{\mathbf{a}}_{k},\tilde{b}_{k})$ that the iterative algorithm produces lies inside the cone of the set of vectors $\{\sum_{i\in I_{jk}}|c_{ji}|(\mathbf{a}_{i}^{T},b_{i})|j\in[m]\}$, which is a subset of the cone of the set of vectors $\{(\mathbf{a}_{i}^{T},b_{i})|i\in I_{k}\}$. Thus, the representative forms an acute angle with every vector of the set $\{(\mathbf{a}_{i}^{T},b_{i})|i\in I_{k}\}$, and thus $x_{ji}\geq 0$.
For every cluster $k\in C_{j+}$ define $I_{jk}^{\prime}=I_{k}\cap I_{j+}^{\prime}$ and
$\tilde{x}_{jk}=\sum_{i\in I_{jk}^{\prime}}x_{ji}.$
Since $x_{ji}\geq 0$, and by the definition of the update step for the output weights, we have
$\sum_{i\in I_{jk}}x_{ji}=1\Rightarrow\tilde{x}_{jk}=\sum_{i\in I_{jk}^{\prime}}x_{ji}\leq 1$
Thus, for every cluster $k\in C_{j+}$, the point $\tilde{x}_{jk}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ lies inside the segment $\mathbf{0},\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$], and thus belongs to $\tilde{P_{j}}$.
For the vertex $\mathbf{u}$, we choose to compare it with the point
$\sum_{k\in C_{j+}}\tilde{x}_{jk}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\in\tilde{P_{j}}$.
We have that
$\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})=\sum_{i\in I_{jk}}c_{ji}(\mathbf{a}_{i}^{T},b_{i})+l_{jk}=|I_{jk}|\frac{\sum_{i\in I_{jk}}c_{ji}(\mathbf{a}_{t}^{T},b_{i})}{|I_{jk}|}+l_{jk}=|I_{jk}|(c_{ji}(\mathbf{a}_{i}^{T},b_{i})+\epsilon_{ji})+l_{jk},$
where $\sum_{j=1}^{m}\|l_{jk}\|^{2}=l_{k}^{2}$ the optimization criterion loss, and $\epsilon_{ji}$ is the error between $c_{ji}(\mathbf{a}_{i}^{T},b_{i})$ and the mean $\frac{\sum_{i\in I_{jk}}c_{ji}(\mathbf{a}_{t}^{T},b_{i})}{|I_{jk}|}$. It is easy to verify that $\epsilon_{ji}$ tends to zero as $\delta_{max}$ tends to 0.
Thus, we have that
$\operatorname{dist}(\mathbf{u},\tilde{P}_{j})$ $\leq\left\|\sum_{i\in I_{j+}^{\prime}}c_{ji}(\mathbf{a}_{i}^{T},b_{i})-\sum_{k\in C_{j+}}\tilde{x}_{jk}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|$
$\leq\sum_{k\in C_{j+}}\left\|\sum_{i\in I_{jk}^{\prime}}c_{ji}(\mathbf{a}_{i}^{T},b_{i})-\tilde{x}_{jk}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
$\leq\sum_{k\in C_{j+}}\left\|\sum_{i\in I_{jk}^{\prime}}\left[c_{ji}(\mathbf{a}_{i}^{T},b_{i})-x_{ji}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right]\right\|+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
$\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}^{\prime}}\left\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})-x_{ji}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
$\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}^{\prime}}\min_{x}\Bigl\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})-x\left(|I_{jk}|(c_{ji}(\mathbf{a}_{i}^{T},b_{i})+\epsilon_{ji})+l_{jk}\right)\Bigr\|+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
$\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}^{\prime}}\min_{x}\Bigl\|(1-x|I_{jk}|)c_{ji}(\mathbf{a}_{i}^{T},b_{i})-xl_{jk}-x|I_{jk}|\epsilon_{ji}\Bigr\|+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
$\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}^{\prime}}\min_{x}\Biggl\{|1-x|I_{jk}||\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|+|x|I_{jk}||\left\|\frac{l_{jk}}{|I_{jk}|}+\epsilon_{ji}\right\|\Biggr\}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
$\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}^{\prime}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\left\|\frac{l_{jk}}{|I_{jk}|}+\epsilon_{ji}\right\|\right\}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
$\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}^{\prime}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|.$
The maximum value of the upper bound occurs when $I_{j+}^{\prime}=I_{j+}$. Thus, we have
$\max_{\mathbf{u}\in V_{P_{j}}}\operatorname{dist}(\mathbf{u},\tilde{P}_{j})\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|.$
To obtain a bound for $\max_{\mathbf{v}\in V_{\tilde{P}_{j}}}\operatorname{dist}(P_{j},\mathbf{v})$, we write $\mathbf{v}=\sum_{k\in C_{j+}^{\prime}}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\in\tilde{P}_{j}$ and choose vertex $\mathbf{u}=\sum_{i\in I_{j+}^{\prime}}c_{ji}(\mathbf{a}_{i}^{T},b_{i})$, with $I_{j+}^{\prime}=\{i\in I_{j+}|i\in I_{k},k\in C_{j+}^{\prime}\}$. For this set we have $\tilde{x}_{jk}=1$, and thus this distance has already been taken into account in the calculation of $\max_{\mathbf{u}\in V_{P_{j}}}\operatorname{dist}(\mathbf{u},\tilde{P}_{j})$.
At last, we have
$H(P_{j},\tilde{P}_{j})\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|.$
Similarly, for $H(Q_{j},\tilde{Q}_{j})$ we have
$H(Q_{j},\tilde{Q}_{j})\\ \leq\sum_{k\in C_{j-}}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}\\ +\sum_{i\in N_{j-}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|.$
Combining, we obtain
$\frac{1}{\rho}\max_{\mathbf{x}\in B}|v_{j}(\mathbf{x})-\tilde{v}_{j}(\mathbf{x})|\leq\sum_{k=1}^{m}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}\\ +\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|.$
Using the fact that $I_{jk}\subseteq I_{k}$ and $N_{min}\leq|I_{jk}|,\forall j,k$ we have
$\frac{1}{\rho}\max_{\mathbf{x}\in B}|v_{j}(\mathbf{x})-\tilde{v}_{j}(\mathbf{x})|\leq\sum_{k=1}^{m}\sum_{i\in I_{k}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{N_{min}}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|.$
We make use of the following inequality, which is a direct consequence of Cauchy-Schwartz Inequality
$\sum_{j=1}^{m}|u_{j}|\leq\sqrt{m}\sqrt{\sum_{j=1}^{m}|u_{j}|^{2}}=\sqrt{m}\|(u_{1},\dots,u_{m})\|.$
We have
$\sum_{j=1}^{m}|c_{ji}|\leq\sqrt{m}\|\mathbf{C}_{:,i}\|,$
$\sum_{j=1}^{m}\|l_{jk}\|\leq\sqrt{m}\sqrt{\sum_{j=1}^{m}\|l_{jk}\|^{2}}=\sqrt{m}\cdot l_{k},$
$\sum_{j=1}^{m}\|\epsilon_{ji}\|\leq\sqrt{m}\sqrt{\sum_{j=1}^{m}\|\epsilon_{ji}\|^{2}}=\sqrt{m}\|\epsilon_{:,i}\|_{F}.$
Using the above inequalities, the fact that $\sum\min\leq\min\sum$, and the fact that $\max\sum\leq\sum\max$ we get
$\frac{1}{\rho}\max_{\mathbf{x}\in B}\|v(\mathbf{x})-\tilde{v}(\mathbf{x})\|_{1}$ $\leq\sum_{k=1}^{m}\sum_{i\in I_{k}}\min\left\{\sum_{j=1}^{m}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\sum_{j=1}^{m}\|l_{jk}\|}{N_{min}}+\sum_{j=1}^{m}\|\epsilon_{ji}\|\right\}$
$+\sum_{j=1}^{m}\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
$\leq\sqrt{m}\sum_{i=1}^{n}\min\left\{\|\mathbf{C}_{:,i}\|\|(\mathbf{a}_{i}^{T},b_{i})\|,\frac{l_{k(i)}}{N_{min}}+\|\epsilon_{:,i}\|_{F}\right\}$
$+\sum_{j=1}^{m}\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|,$
as desired. ∎